Mir Status Report 13 April 1997

NASA MIR PROGRAM
Mir 23 / NASA 4 Status
Date: 13 April 1997
Mission Day: Mir 23 / NASA 4 MD 63 / 92

Mir Station System/Subsystem Activities Mir IFM Repair (entire crew) Rest (entire crew)

SMP/Medical: per day according to the first day of the exercise Cyclogram (treadmill and expanders). A radiogram from the Russian medical group has been prepared for uplink today or tomorrow, recommending that the commander and Jerry return to two hours of exercise per day in preparation for the EVA.

Private Family Conference
Successful 11 minute audio PFC between Jerry and Kathryn.

Mir Systems
Friday and Saturday, the Elektron system continued to shut down automatically after runs varying in duration between 3 and 20 minutes. Since last night, the system has run continuously for 22 hours without a problem (as of 1445 DMT). The crew was instructed to stop using air from the Progress, since the Elektron is now producing sufficient oxygen.

The Vosdukh system is removing CO2. The ppCO2 is now about 4.8 mmHg. The system is operating in an `off-nominal' mode. The lack of cooling (due to the VGK shutdown) is causing the systems water vapor removal to be less efficient, resulting in more water than usual being vented overboard with the CO2.

Leak checks continue on the VGK system in Kvant. There is some concern that the leak may be in a hydroconnector at a point with little or no access. More panels will be opened for examination tomorrow.

The Urine Processing System is also operating in an off-nominal mode. A Radiogram has been sent up to the crew with instructions on how to over-ride or work-around a problem with a micro-pump in the system. Urine is being transferred to one of the tanks on the Progress.

Environmental Parameters (Information from 12 April)

Parameter Min. Max. Nominal Range       Pressure (mmHg)  746 755 660 - 850
ppO2 (mmHg) 159 162 140 -200
ppCO2 (mmHg) 4.2 4.7 up to 8.0
ppH20 (mmHg) 12.3 13.1 8 - 14
Humidity (percentage) 40 42 30 - 75
Temperature (C) (BB) 28.5 28.8 18 - 28
Temp (C) transfer module 21.7 22.2 18 - 28
Temp (C) work area Krystall 25.7 25.8 18 - 28
Temp (C) sleep area Krystall 27.0 27.3 18 - 28
Temp (C) Spektr 25.5 25.8 18 - 28
Temp (C) Kvant -- 18 - 28


Planning

MIR 23/NASA 4 MD 64/93
KOB-2 Loop Leak Repair (entire crew) QUELD Sample E-13 (NASA 4)
DOZA - TLD Processing (NASA
4)

  • submit to reddit


Loading



Humans to Mars Summit 2019
Military Space Situational Awareness April 1 - April 2, 2019, London, UK
MilSatCom USA conference will return to Washington from the 26th – 27th June 2019
Space Access Conference 2019
Mission To Methonē by Les Johnson - Baen Books




Join our mailing list




Monthly Archives

About this Entry

This page contains a single entry by Keith Cowing published on April 13, 1997 12:37 PM.

Mir Status Report 12 April 1997 was the previous entry in this blog.

Mir Status Report 14 April 1997 is the next entry in this blog.

Find recent content on the main index or look in the archives to find all content.